40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10802     Entry time: Tue Dec 16 00:20:06 2014
Author: diego 
Type: Update 
Category: Optical Levers 
Subject: BS & PRM OL realignment 

[Rana, Diego]

We manually realigned the BS and PRM optical levers on the optical table.

ELOG V3.1.3-